# Theoretical Detailed Analyses for DC readout and a Fabri-Pérot gravitational-wave detector

### v1

@@ VII DC readout scheme for a Fabri-Pérot Graviational-Wave Detector
$D_{d}^{\dagger}\hat{b}_{n}(\omega_{0}\pm\Omega)D_{d}=D_{d}^{\dagger}\hat{b}_{n\pm}(\Omega)D_{d}$ (215)
@@
We note that the additional classical part $\mathfrak{R}$}} defined by Eq. (LABEL:eq:FrakR-def) in $\mathfrak{B}$}} of Eq. (213) does not contribute to the expectation value (219) due to the reality condition of the Fourier transformation $H(\Omega,L)$ of the gravitational-wave signal. Here, we note that the expectation value (219) is maximized when the offset $\theta=\pi/4$, while we do not measure $H(\Omega,L)$ when the complete dark port $\theta=0$. In the case of $\theta=0$, the gravitational-wave signal $H(\Omega,L)$ is included in the term of $O((I_{0})^{0})$. This is the well-known fact that the complete dark-port condition $\theta=0$ is meaningless in the DC-readout scheme. The factor $\sin(2\theta)$ in Eq. (219) comes from the fact that ${{\mathfrak{A}}}(\omega_{0}\pm\Omega)$ in Eq. (213) depends on $\cos\theta$ and $\mathfrak{B}$}} in Eq. (214) is proportional to $\sin\theta$. Thus, we concludes that $\theta=\pi/4$ maximize the expectation value (219).
@@
Next, we consider the stationary noise spectral density (48). From the explicit expression (215) of the output operator $D_{d}^{\dagger}\hat{b}_{n\pm}(\Omega)D_{d}$, the amplitude- and phase-quadrature $D_{d}^{\dagger}\hat{b}_{n1}(\Omega)D_{d}$ and $D_{d}^{\dagger}\hat{b}_{n2}(\Omega)D_{d}$ defined by Eqs. (45) and (46), respectively, are given by
$D_{d}^{\dagger}\hat{b}_{n1}(\Omega)D_{d}$
@@
$D_{d}^{\dagger}\hat{b}_{n2}(\Omega)D_{d}$
@@
From the expressions (LABEL:eq:B1-noise-final) and (LABEL:eq:B2-noise-final) of the amplitude- and phase-quadrature, the operator $D_{d}^{\dagger}\hat{b}_{n\Theta}D_{d}$ defined by Eq. (47) is given by
$D_{d}^{\dagger}\hat{b}_{n\Theta}(\Omega)D_{d}$
@@
$D_{d}^{\dagger}\hat{b}_{n\Theta}^{\dagger}(\Omega)D_{d}$
@@
Through the operators $D_{d}^{\dagger}\hat{b}_{n\Theta}(\Omega)D_{d}$ above, the straightforward calculations for the right-hand side of Eq. (48) yields
@@
Figure 6: The square root of the noise spectral density $\sqrt{S_{H}}$ for the $m_{EM}=m_{ITM}$ and the optimal offset $\theta=\pi/4$ case. (EQ) indicates $m_{EM}=m_{ITM}$ case. The noise spectral density $S_{H}$ is defined by Eq. (234). $\epsilon$ is the incomplete parameter discussed Sec. VIII.
Figure 7: The square root of the noise spectral density $\sqrt{S_{H}}/{{\mathfrak{R}}}$ with equal mass $(EQ)$ case where $m_{EM}=m_{ITM}$ and the offset is $\theta=\frac{\pi}{4}$. The behavior of this ratio is similar to Kimble’s signal-referred noise spectral density $\sqrt{S_{K}}$.
@@
in Fig. 8 to clarify the difference between the derived noise-spectral density $S_{H}$ and the original Kimble’s noise spectral density. This ratio is approximately unity except for the high frequency range $\sim 20$kHz. At the high frequency range $\sim 20$kHz, the noise spectral density $\sqrt{2S_{H}^{(EQ)}}/{{\mathfrak{R}}}$ is 20% smaller than the noise spectral density $S_{(K)}^{(K)}$ defined by Eq. (212). We also observe a small power dependence around 10 Hz, indicating roughly 1% differences from $S_{(K)}^{(K)}$. Since the expectation value $|\langle\hat{s}_{{{\mathcal{N}}}_{b}}(\Omega)\rangle|^{2}$ given by Eq. (219) is the largest when $\theta=\pi/4$, we show Figures 6, 7 and 8 with $\theta=\pi/4$.
Figure 8: The square root ratio of the noise spectral density $S_{H}^{(EQ)}/{{\mathfrak{R}}}^{2}$ and $S_{(K)}^{(K)}$ where $S_{(K)}^{(K)}$ is defined by Eq. (212). This ratio is approximately unity except for the high frequency range $\sim 20$kHz. At the high frequency range $\sim 20$kHz, the noise spectral density $\sqrt{2S_{H}^{(EQ)}}/{{\mathfrak{R}}}$ is 20% smaller than the noise spectral density $S_{(K)}^{(K)}$. We also observe a small power dependence around 10 Hz, indicating roughly 1% differences from $S_{(K)}^{(K)}$.
@@
The factor $\mathfrak{R}$}} in the signal referred noise spectral density $S_{H}(\Omega)$ comes from (214). Equation (214) indicates that the classical carrier field is modified by $\mathfrak{R}$}}. In DC readout schemes, we use the classical carrier field as the reference to measure gravitational-wave signals. Due to this modification of the classical carrier field appears to the signal referred noise spectral density $S_{H}(\Omega)$, although it does not appear in the expectation value (219) of the output signal due to the reality condition of the gravitational-wave signals $H(\Omega)=H^{*}(-\Omega)$. Furthermore, we also note that the existence of the factor $\mathfrak{R}$}} leads to the signal-referred noise spectral density in which the shot noise is not reduced even if the injected power $I_{0}$ is increased. This contradicts the consensus in the community of gravitational-wave experiments.
@@
Originally, the additional classical carrier field $\mathfrak{R}$}} comes from the solutions (172) and (174) to the Heisenberg equations as the terms proportional to $2\pi\delta(\Omega)$. Furthermore, these terms come from the radiation pressure forces (162)–(165) as the classical constant force in the time domain. This classical constant force in the time domain appears in the Heisenberg equations of motion (141), (142), (145), and (146). In both classical and quantum mechanics of a forced harmonic oscillator, the constant force in the harmonic oscillator leads to changes in the equilibrium points of the oscillator. In actual experiments, the interferometers should be operated at these equilibrium points of mirrors. Therefore, to take into account the effects of the constant forces, we have to change the tuning condition (161) so that the laser is tuned at the equilibrium points of mirrors. Therefore, we have to consider the change of the tuning condition (161) in the next section VIII. To do this, we have to reconsider the Heisenberg equations for mirrors in Sec. V.
@@ VIII.1 Complete equilibrium tuning
Now, we consider the modification of the tuning points from the tuning conditions (161) and (177). First, we consider the modification of the tuning condition (161). To do this, we note that the cavity propagation conditions (78) and (79) yields that the retarded effects in the electric field operators are determined by
@@
where $\hat{X}_{x}^{(fluc)}:=\hat{X}^{(fluc)}_{XEM}-\hat{X}_{XITM}^{(fluc)}$ and $\hat{X}_{y}^{(fluc)}:=\hat{X}^{(fluc)}_{YEM}-\hat{X}_{YITM}^{(fluc)}$. Here, $\tau^{(o)}$ is regarded as the operation point of the cavity length of the Fabri-Pérot interferometer. Equations (LABEL:eq:retarded-effect-cavity-X-renorm) and (LABEL:eq:retarded-effect-cavity-Y-renom) indicate that we may replace
@@
As in the cases of Eqs. (LABEL:eq:retarded-effect-cavity-X-renorm) and (LABEL:eq:retarded-effect-cavity-Y-renom), $\tau_{x}^{(o)^{\prime}}$ and $\tau_{y}^{(o)^{\prime}}$ is regarded as the operation points of the length between BS and ITMs in the Fabri-Pérot interferometer. Equation (255) and (256) indicate that we may replace
@@
Since all $\hat{X}_{x,y}$, $\hat{X}_{XITM}$, and $\hat{X}_{YITM}$ are replaced by $\hat{X}_{x,y}^{(fluc)}$, $\hat{X}_{XITM}^{(fluc)}$, and $\hat{X}_{YITM}^{(fluc)}$, respectively, their Fourier transformations $\hat{Z}_{x,y}$, $\hat{Z}_{XITM}$, and $\hat{Z}_{YITM}$ are also replaced to $\hat{Z}_{x,y}^{(fluc)}$, $\hat{Z}_{XITM}^{(fluc)}$, and $\hat{Z}_{YITM}^{(fluc)}$. These $\hat{Z}_{x,y}^{(fluc)}$, $\hat{Z}_{XITM}^{(fluc)}$, and $\hat{Z}_{YITM}^{(fluc)}$ does not includes the classical term which proportional to $2\pi\delta(\Omega)$. Therefore, the resulting input-output relation corresponding to Eq. (189) is given by the expression of Eq. (189) with ${{\mathfrak{R}}}=0$. In this expression of the input-output relation, there is no additional term in the classical carrier fields $\mathfrak{R}$}}, we obtain the noise spectral density $S_{H}(\Omega)=S_{(K)}(\Omega)/2$ instead of Eq. (234). This recover the conventional noise-spectral density except for the fact that we have to use the definitions (187) and (188) of the variables $h_{SQL}$ and $\kappa$ with the replacement $\tau\rightarrow\tau^{(o)}$, respectively. The difference between the resulting noise spectral density $S_{H}(\Omega)=S_{(K)}(\Omega)/2$ and Kimble’s noise spectral density $S_{(K)}^{(K)}$ [8] are depicted in Fig. 5 except for the factor $1/2$ in $S_{H}(\Omega)$ as discussed in Sec. VII.
@@ VIII.2 Incomplete equilibrium tuning
If we achieve the above complete equilibrium tuning in experiments, we can realize Kimble’s noise spectral density [8] with the modifications of $\kappa$ and $h_{SQL}$ as an idealized case. However, in some experiments, we might not be able to accomplish the above complete equilibrium tuning for some reasons. In this subsection, we consider the case where we cannot realize the above complete equilibrium tuning and how this incompleteness appears in the noise spectral densities through our consideration within this paper.
@@
through ${{\mathscr{D}}}_{ITM}$, $\hat{X}_{XITM}^{(fluc)}(t)$, and $\hat{X}_{YITM}^{(fluc)}(t)$ defined by Eqs. (240), (245), and (246), respectively. Here, $\epsilon$ is a phenomenological parameter within $[0,1]$.
As in the case of the complete equilibrium tuning, we regard $\tau_{x}^{\epsilon(o)^{\prime}}$ and $\tau_{y}^{\epsilon(o)^{\prime}}$ as the operation points of the length between BS and ITMs in the Fabri-Pérot interferometer. Equations (270) and (272) indicate that we may replace
@@
$\tau_{y}^{\prime}\rightarrow\tau_{y}^{\epsilon(o)^{\prime}},$ (275)
@@
$\hat{X}_{YITM}(t)\rightarrow\hat{X}_{YITM}^{\epsilon(o)}(t)$ (277)
@@
As in the case of the complete equilibrium tuning, we regard $\tau^{\epsilon(o)}$ as the operation point of the cavity length between ITMs and EMs in the Fabri-Pérot interferometer. Equations (278) and (LABEL:eq:L-incomplete-operation-point-y) indicate that we may replace
@@
Figure 9: The schematic picture of the incomplete equilibrium tuning. The phenomenological parameter $\epsilon$ represents the deviations of the tuning point from the equilibrium point of the pendulum for the mirrors. $\epsilon=0$ corresponds to the complete equilibrium tuning. Even if we keep the mirrors at the operation point $\epsilon\neq 0$, the mirrors still feel acceleration and try to move. These accelerations for each mirror are included in the feedback electric current.
@@
Note that $\hat{X}_{XITM}={{\mathscr{D}}}_{ITM}$, $\hat{X}_{YITM}={{\mathscr{D}}}_{ITM}$, $\hat{X}_{XEM}={{\mathscr{D}}}_{EM}$, and $\hat{X}_{YEM}={{\mathscr{D}}}_{EM}$ are equilibrium points of the pendulum for mirrors, as mentioned in Sec. VIII.1. Even if we make the mirrors stay at the operation point so that $\epsilon\neq 0$, the mirrors feel their accelerations due to the classical radiation pressure and try to move with the frequency $\omega_{p}$ of the pendulum. In this case, these accelerations are measured through the feedback control system, and its signal is included in the feedback electric current. In this case, the noise in the feedback electric current may deviate from the noise at the photodetection. For this reason, we call the tuning conditions $\epsilon\neq 0$ as “incomplete equilibrium tuning.” The schematic picture of the incomplete equilibrium tuning is depicted in Fig. 9.
Figure 10: The square-root of the explicit signal referred noise spectral density $\sqrt{S_{(H)}^{EQ}(\Omega,\epsilon,\theta=0)}$ for each $\epsilon$ is depicted in the range1Hz to 20 kHz with different values of the laser power $I_{0}$. $S_{(H)}^{(EQ)}$ is given by Eq. (293) with $m=m_{EM}=m_{ITM}$.
@@
In Fig. 10, the left-top panel shows the conventional Kimble’s noise spectral density with the modification of $\kappa$ and $h_{SQL}$. However, if the incompleteness $\epsilon$ increases, the deviation from Kimble’s noise spectral density can be seen. After all, in the high-frequency region, the noise spectral density shows that the shot noise does not decrease due to the increase of the incident laser power $I_{0}$. This is merely due to the contribution of the additional classical carrier field $\epsilon{{\mathfrak{R}}}$, which may be dominant in the incomplete equilibrium tuning.
Furthermore, the minimum at Eq. (297) of the noise spectral density can be seen as a dip around the frequency (297) in Fig. 10. This dip shows a violation of the so-called “standard quantum limit” which is the envelope of the left-top figure in Fig. 10. However, as discussed in Sec. V.1, this effect is not the violation of the conventional arguments of the Heisenberg uncertainty principle which arise from the non-commutation of the position operator $\hat{X}$ and the momentum operator $\hat{P}$, i.e., $\hat{X},\hat{P}]=i\hbar$. This is because the quantum mirrors’ initial conditions, which have the information $\hat{X},\hat{P}]=i\hbar$, do not affect our consideration as discussed in Sec. V.1. Therefore, we have to say that this dip has nothing to do with the arguments of the violation of the conventional arguments of the Heisenberg uncertainty principle in Ref. [12]. In the gravitational-wave community, it is well-known that the standard quantum limit can be violated through appropriate changes of the interferometer configurations. Therefore, we may regard the dip in Fig. 10 as one of them.
Figure 11: The square-root of the explicit signal-referred noise spectral density $\sqrt{S_{(H)}^{EQ}(\Omega,\epsilon,\theta=\pi/4)}$ for each $\epsilon$ is depicted in the range 1Hz to 20 kHz with different values of the laser power $I_{0}$. $S_{(H)}^{(EQ)}$ is given by Eq. (298).

### v2

@@ VII DC readout scheme for a Fabry-Pérot Graviational-Wave Detector
${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n}(\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}={{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n\pm}(\Omega){{\mathfrak{D}}}_{d}$ (215)
@@
We note that the additional classical part $\mathfrak{R}$}} defined by Eq. (LABEL:eq:FrakR-def) in $\mathfrak{B}$}} of Eq. (213) does not contribute to the expectation value (219) due to the reality condition of the Fourier transformation $H(\Omega,L)$ of the gravitational-wave signal. Here, we note that the expectation value (219) is maximized when the offset $\theta=\pi/4$, while we do not measure $H(\Omega,L)$ when the complete dark port $\theta=0$. In the case of $\theta=0$, the gravitational-wave signal $H(\Omega,L)$ is included in the term of $O((I_{0})^{0})$. It is well-known that the complete dark-port condition $\theta=0$ is meaningless in the DC-readout scheme. The factor $\sin(2\theta)$ in Eq. (219) comes from the fact that ${{\mathfrak{A}}}(\omega_{0}\pm\Omega)$ in Eq. (213) depends on $\cos\theta$ and $\mathfrak{B}$}} in Eq. (214) is proportional to $\sin\theta$. Thus, we conclude that $\theta=\pi/4$ maximize the expectation value (219).
@@
Next, we consider the stationary noise spectral density (48). From the explicit expression (215) of the output operator ${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n\pm}(\Omega){{\mathfrak{D}}}_{d}$, the amplitude- and phase-quadrature ${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n1}(\Omega){{\mathfrak{D}}}_{d}$ and ${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n2}(\Omega){{\mathfrak{D}}}_{d}$ defined by Eqs. (45) and (46), respectively, are given by
${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n1}(\Omega){{\mathfrak{D}}}_{d}$
@@
${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n2}(\Omega){{\mathfrak{D}}}_{d}$
@@
From the expressions (LABEL:eq:B1-noise-final) and (LABEL:eq:B2-noise-final) of the amplitude- and phase-quadrature, the operator ${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n\Theta}{{\mathfrak{D}}}_{d}$ defined by Eq. (47) is given by
${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n\Theta}(\Omega){{\mathfrak{D}}}_{d}$
@@
${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n\Theta}^{\dagger}(\Omega){{\mathfrak{D}}}_{d}$
@@
Through the operators ${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{n\Theta}(\Omega){{\mathfrak{D}}}_{d}$ above, the straightforward calculations for the right-hand side of Eq. (48) yields
@@
Figure 6: The square root of the noise spectral density $\sqrt{S_{H}}$ for the $m_{EM}=m_{ITM}$ and the optimal offset $\theta=\pi/4$ case. (EQ) indicates $m_{EM}=m_{ITM}$ case. The noise spectral density $S_{H}$ is defined by Eq. (234). $\epsilon$ is the incomplete parameter discussed Sec. VIII, while $\epsilon=0$ corresponds to the fact that the tuning point is the equilibrium point of the pendulum and $\epsilon=1$ corresponds to the tuning without taking into account of the displacement of the equilibrium point of the pendulum due to the classical radiation pressure from the laser. (See Sec. VIII.)
Figure 7: The square root ratio of the noise spectral density $S_{H}^{(EQ)}/{{\mathfrak{R}}}^{2}$ and $S_{(K)}^{(K)}$ where $S_{(K)}^{(K)}$ is defined by Eq. (212). This ratio is approximately unity except in the high-frequency range $\sim 20$ kHz. At the high frequency range $\sim 20$kHz, the noise spectral density $\sqrt{2S_{H}^{(EQ)}}/{{\mathfrak{R}}}$ is 20% smaller than the noise spectral density $S_{(K)}^{(K)}$. We also observe a small power dependence around 10 Hz, indicating roughly 1% differences from $S_{(K)}^{(K)}$.
@@
in Fig. 7 to clarify the difference between the derived noise-spectral density $S_{H}$ and the original Kimble’s noise spectral density. This ratio is approximately unity except for the high-frequency range $\sim 20$kHz. At the high frequency range $\sim 20$kHz, the noise spectral density $\sqrt{2S_{H}^{(EQ)}}/{{\mathfrak{R}}}$ is 20% smaller than the noise spectral density $S_{(K)}^{(K)}$ defined by Eq. (212). We also observe a small power dependence around 10 Hz, indicating roughly 1% differences from $S_{(K)}^{(K)}$. Since the expectation value $|\langle\hat{s}_{{{\mathcal{N}}}_{b}}(\Omega)\rangle|^{2}$ given by Eq. (219) is the largest when $\theta=\pi/4$, we show Figures 6, 7, and 8 with $\theta=\pi/4$.
Figure 8: The square root of the noise spectral density $\sqrt{S_{H}}/{{\mathfrak{R}}}$ with equal mass $(EQ)$ case where $m_{EM}=m_{ITM}$ and the offset is $\theta=\pi/4$. The behavior of this ratio is similar to Kimble’s signal-referred noise spectral density $\sqrt{S_{K}}$.
@@
The factor $\mathfrak{R}$}} in the signal referred noise spectral density $S_{H}(\Omega)$ comes from (214). Equation (214) indicates that the classical carrier field is modified by $\mathfrak{R}$}}. In DC readout schemes, we use the classical carrier field as the reference to measure gravitational-wave signals. Due to this modification of the classical carrier field appears to the signal referred noise spectral density $S_{H}(\Omega)$, although it does not appear in the expectation value (219) of the output signal due to the reality condition of the gravitational-wave signals $H(\Omega)=H^{*}(-\Omega)$. Furthermore, we also note that the existence of the factor $\mathfrak{R}$}} leads to the signal-referred noise spectral density in which the shot-noise contribution is not reduced even if the injected power $I_{0}$ is increased. This contradicts the consensus in the community of gravitational-wave experiments.
@@
Originally, the additional classical carrier field $\mathfrak{R}$}} comes from the solutions (172) and (LABEL:eq:Eq-of-motion-XYITM-com-tuned-wobifreq-sol-sum-approx) to the Heisenberg equations as the terms proportional to $2\pi\delta(\Omega)$. Furthermore, these terms come from the radiation pressure forces (162)–(165) as the classical constant force in the time domain. This classical constant force in the time domain appears in the Heisenberg equations of motion (141), (142), (145), and (146). In both classical and quantum mechanics of a forced harmonic oscillator, the constant force in the harmonic oscillator leads to changes in the equilibrium points of the oscillator. In actual experiments, the interferometers should be operated at the mirrors’ equilibrium positions. Therefore, to take into account the effects of the constant forces, we have to change the tuning condition (161) so that the laser is tuned at the equilibrium points of mirrors. Therefore, we have to consider the change of the tuning condition (161) in the next section VIII. To do this, we have to reconsider the Heisenberg equations for mirrors in Sec. V.
@@ VIII.1 Complete equilibrium tuning
Now, we consider the modification of the tuning points from the tuning conditions (161) and (177). First, we consider the modification of the tuning condition (161). To do this, we note that the cavity propagation conditions (78) and (79) yield that the retarded effects in the electric field operators are determined by
@@
where $\hat{X}_{x}^{(fluc)}:=\hat{X}^{(fluc)}_{XEM}-\hat{X}_{XITM}^{(fluc)}$ and $\hat{X}_{y}^{(fluc)}:=\hat{X}^{(fluc)}_{YEM}-\hat{X}_{YITM}^{(fluc)}$. Here, $\tau^{(o)}$ is regarded as the operation point of the cavity length of the Fabry-Pérot interferometer. Equations (LABEL:eq:retarded-effect-cavity-X-renorm) and (LABEL:eq:retarded-effect-cavity-Y-renom) indicate that we may replace
@@
As in the cases of Eqs. (LABEL:eq:retarded-effect-cavity-X-renorm) and (LABEL:eq:retarded-effect-cavity-Y-renom), $\tau_{x}^{(o)^{\prime}}$ and $\tau_{y}^{(o)^{\prime}}$ is regarded as the operation points of the length between BS and ITMs in the Fabry-Pérot interferometer. Equation (255) and (256) indicate that we may replace
@@
Since all $\hat{X}_{x,y}$, $\hat{X}_{XITM}$, and $\hat{X}_{YITM}$ are replaced by $\hat{X}_{x,y}^{(fluc)}$, $\hat{X}_{XITM}^{(fluc)}$, and $\hat{X}_{YITM}^{(fluc)}$, respectively, their Fourier transformations $\hat{Z}_{x,y}$, $\hat{Z}_{XITM}$, and $\hat{Z}_{YITM}$ are also replaced to $\hat{Z}_{x,y}^{(fluc)}$, $\hat{Z}_{XITM}^{(fluc)}$, and $\hat{Z}_{YITM}^{(fluc)}$. These $\hat{Z}_{x,y}^{(fluc)}$, $\hat{Z}_{XITM}^{(fluc)}$, and $\hat{Z}_{YITM}^{(fluc)}$ does not includes the classical term which proportional to $2\pi\delta(\Omega)$. Therefore, the resulting input-output relation corresponding to Eq. (189) is given by the expression of Eq. (189) with ${{\mathfrak{R}}}=0$. In this expression of the input-output relation, there is no additional term in the classical carrier fields $\mathfrak{R}$}}, we obtain the noise spectral density $S_{H}(\Omega)=S_{(K)}(\Omega)/2$ instead of Eq. (234). This recovers the conventional noise-spectral density except for the fact that we have to use the definitions (187) and (188) of the variables $h_{SQL}$ and $\kappa$ with the replacement $\tau\rightarrow\tau^{(o)}$, respectively. The difference between the resulting noise spectral density $S_{H}(\Omega)=S_{(K)}(\Omega)/2$ and Kimble’s noise spectral density $S_{(K)}^{(K)}$ [10] are depicted in Fig. 5 except for the factor $1/2$ in $S_{H}(\Omega)$ as discussed in Sec. VII.
@@ VIII.2 Incomplete equilibrium tuning
If we achieve the above complete equilibrium tuning in experiments, we can realize Kimble’s noise spectral density [10] with the modifications of $\kappa$ and $h_{SQL}$ as an idealized case. However, in some experiments, we might not be able to achieve complete equilibrium tuning for various reasons. In this subsection, we consider the case where we cannot realize the above complete equilibrium tuning and how this incompleteness appears in the noise spectral densities through our consideration within this paper.
@@
through ${{\mathscr{D}}}_{ITM}$, $\hat{X}_{XITM}^{(fluc)}(t)$, and $\hat{X}_{YITM}^{(fluc)}(t)$ defined by Eqs. (240), (245), and (246), respectively. Here, $\epsilon$ is a parameter of incompleteness of the tuning within $[0,1]$. We choose $\epsilon$ so that $\epsilon=0$ corresponds to the tuning point, which is the equilibrium point of the pendulum, and $\epsilon=1$ corresponds to the tuning without taking into account the displacement of the equilibrium point of the pendulum due to the classical radiation pressure from the laser.
As in the case of the complete equilibrium tuning, we regard $\tau_{x}^{\epsilon(o)^{\prime}}$ and $\tau_{y}^{\epsilon(o)^{\prime}}$ as the operation points of the length between BS and ITMs in the Fabry-Pérot interferometer. Equations (270) and (272) indicate that we may replace
@@
$\tau_{y}^{\prime}\rightarrow\tau_{y}^{\epsilon(o)^{\prime}},$
@@
$\hat{X}_{YITM}(t)\rightarrow\hat{X}_{YITM}^{\epsilon(o)}(t)$
@@
As in the case of the complete equilibrium tuning, we regard $\tau^{\epsilon(o)}$ as the operation point of the cavity length between ITMs and EMs in the Fabry-Pérot interferometer. Equations (LABEL:eq:L-incomplete-operation-point-x) and (LABEL:eq:L-incomplete-operation-point-y) indicate that we may replace
@@
Figure 9: The schematic picture of the incomplete equilibrium tuning. The incomplete parameter $\epsilon$ represents the deviations of the tuning point from the equilibrium point of the pendulum for the mirrors. $\epsilon=0$ corresponds to the complete equilibrium tuning. Even if we keep the mirrors at the operation point $\epsilon\neq 0$, the mirrors still feel acceleration and try to move. These accelerations for each mirror are included in the feedback electric current.
@@
Note that $\hat{X}_{XITM}={{\mathscr{D}}}_{ITM}$, $\hat{X}_{YITM}={{\mathscr{D}}}_{ITM}$, $\hat{X}_{XEM}={{\mathscr{D}}}_{EM}$, and $\hat{X}_{YEM}={{\mathscr{D}}}_{EM}$ are equilibrium points of the pendulum for mirrors, as mentioned in Sec. VIII.1. Even if we make the mirrors stay at the operation point so that $\epsilon\neq 0$, the mirrors feel their accelerations due to the classical radiation pressure and try to move with the frequency $\omega_{p}$ of the pendulum. In this case, these accelerations are measured by the feedback control system, and their signals are included in the feedback electric current. In this case, the noise in the feedback electric current may differ from that at the photodetection. For this reason, we call the tuning conditions $\epsilon\neq 0$ as “incomplete equilibrium tuning.” The schematic picture of the incomplete equilibrium tuning is depicted in Fig. 9.
Figure 10: The square root of the explicit signal referred noise spectral density $\sqrt{S_{(H)}^{EQ}(\Omega,\epsilon,\theta=0)}$ for each $\epsilon$ is depicted in the range 1Hz to 20 kHz with different values of the laser power $I_{0}$. $S_{(H)}^{(EQ)}$ is given by Eq. (293) with $m=m_{EM}=m_{ITM}$.
@@
In Fig. 10, the left-top panel shows the conventional Kimble’s noise spectral density with the modification of $\kappa$ and $h_{SQL}$. However, if the incompleteness $\epsilon$ increases, the deviation from Kimble’s noise spectral density can be seen. After all, in the high-frequency region, the noise spectral density shows that the shot-noise contribution in the signal-referred noise spectral density does not decrease due to the increase of the incident laser power $I_{0}$. This is merely due to the contribution of the additional classical carrier field $\epsilon{{\mathfrak{R}}}$, which may be dominant in the incomplete equilibrium tuning.
Furthermore, the minimum at Eq. (297) of the noise spectral density can be seen as a dip around the frequency (297) in Fig. 10. This dip shows a violation of the so-called “standard quantum limit” which is the envelope of the left-top figure in Fig. 10. However, as discussed in Sec. V.1, this effect is not the violation of the conventional arguments of the Heisenberg uncertainty principle which arise from the non-commutation of the position operator $\hat{X}$ and the momentum operator $\hat{P}$, i.e., $\hat{X},\hat{P}]=i\hbar$. This is because the quantum mirrors’ initial conditions, which have the information $\hat{X},\hat{P}]=i\hbar$, do not affect our consideration as discussed in Sec. V.1. Therefore, we have to say that this dip has nothing to do with the arguments of the violation of the conventional arguments of the Heisenberg uncertainty principle in Ref. [14]. In the gravitational-wave community, it is well known that the standard quantum limit can be surpassed using squeezed laser states. This indicates that violating the “standard quantum limit” does not imply violating a physical principle. Therefore, we may regard the dip in Fig. 10 as not a violation of the physical principle.
Figure 11: The square root of the explicit signal-referred noise spectral density $\sqrt{S_{(H)}^{EQ}(\Omega,\epsilon,\theta=\pi/4)}$ for each $\epsilon$ is depicted in the range 1Hz to 20 kHz with different values of the laser power $I_{0}$. $S_{(H)}^{(EQ)}$ is given by Eq. (298).